# Generalizing Scaling Laws for Dense and Sparse Large Language Models

### v1

@@ I Introduction
In recent years, transformer architectures [1] have revolutionized the deep learning approach. These architectures are now the foundation for the majority of popular large language models (LLMs). The reason for the widespread adoption of these architectures is the attention mechanism, which captures the deep contextual relationship in the text data. Because of the large size of LLM models and the large amount of data needed for training, however, the overall training of LLMs becomes time-consuming and expensive. As a result, determining the optimal allocation of resources, model size, and training data and estimating the training performance before conducting the actual training are crucial and have become an important topic in artificial intelligence research. Several approaches have been explored to study the scaling behavior of transformer models. The scaling laws allow researchers and developers to determine the optimal allocation of compute, to select optimal model size and data volume for training foundation models efficiently, and to estimate the performance prior to training. Until now, however, nearly all the approaches [2, 3, 4, 5, 6, 7, 8] have been architecture specific (dense or sparse). In this work, we generalize the existing empirical scaling laws and propose a generalized scaling law for both dense and sparse models that enables accurate performance prediction and optimal resource allocation across different architectures.
@@
Scaling laws—empirical relationships often expressed as power laws—describe a model’s performance based on allocated resources. These laws help determine the optimal model size and data volume for a given compute budget to achieve optimal pretraining performance. Several such empirical scaling laws based on different LLM architectures have been proposed in recent years.[2, 3, 4, 5, 29, 30, 7, 8, 31] Others are based on training precision [30], distillation [32], and quantization [29]. Several studies also have been done on vision [4], audio [33, 34], and mixed-modal language models [35]. This interest in scaling laws for different domains makes them pivotal in deep neural network research.
@@ II-A Dense Scaling Laws
Traditional transformer models are dense in nature. Kaplan et al. [2] studied the behavior of LLMs with a series of experiments and different hyperparameters: model size, model shape, total compute, data volume, and batch size. The authors observed that performance or model loss during pretraining follows a power law relationship with the total number of parameters, the total number of training tokens, and the total compute budget. They established that model loss depends strongly on scale, which consists of the number of model parameters $N$, the dataset size (number of tokens) $D$, and the amount of compute $C$ used for training. It also depends mildly on other architectural hyperparameters such as model shape: depth and width. For dense models, the compute $C$ (training FLOPs) required to train a transformer model with $N$ parameters and $D$ tokens is $C=6ND$ [2] because approximately 6 floating point operations are needed for every parameter in the model and every training sample. They suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times. They represented training loss as a function of number of parameters and data size and proposed the following empirical scaling law:
$L(N,D)=\left[\left(\frac{N_{C}}{N}\right)^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{C}}{D}\right]^{\alpha_{D}}.$ (1)
@@
Using three different approaches—fixed model size with varying numbers of training tokens, IsoFLOP profiling, and parametric function fitting—Hoffmann et al.  [3] extended the work in [2] and concluded that model size and total number of tokens should be increased in proportion. In other words, if the total number of parameters or model size is doubled, data volume or total number of tokens should be doubled as well. They represented training loss as a combination of irreducible loss or entropy $e$, a parameter-dependent term or the loss incurred by a model of fixed parameter size, and a data size term or the loss induced by a fixed number of training steps. Based on their observations, the authors proposed the following scaling law:
@@ II-B1 Pruned Transformer Models
$L(N,D,S)=e+\left(a_{S}(1-S)^{b_{S}}+c_{S}\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (4)
@@ II-B2 MoE Models
$L(N,D,S)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}+\frac{c}{(1-S)^{\lambda}}+\frac{d}{(1-S)^{\delta}N^{\gamma}}.$ (5)
@@ III Generalizing Scaling Laws
Where $0\leq S<1$ because the active parameters in a model cannot be zero. Therefore $\frac{N}{1-S}$ is the total number of parameters. In the parameter-dependent second term in Equation 2, we replace $N$ with $\frac{N}{1-S}$; the parameter-dependent term thus becomes $\frac{a}{\left({\frac{N}{1-S}}\right)^{\alpha}}$, which is simplified as $\frac{a(1-S)^{\alpha}}{N^{\alpha}}$. Similar to the approach in [4] we introduce an upper-bound constant sparsity factor $c$, which is caused by the sparsity; but this factor, which contributes to the loss, depends on the sparsity $S$. So we define the parameter-dependent term as $\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}$. For sparsity $S=0$ or dense models, this term is the same as the original parameter-dependent term in Equation 2.
We have noticed that different approaches adjusted the entropy $e$ in different ways, as shown in Tables II, IV, and V. The value of $e$ is impacted by the sparsity. We adjust the value of $e$ by multiplying it with $(1-S)^{\gamma}$. Accumulating these changes, we propose the following scaling law:
$L(N,D,S)=e(1-S)^{\gamma}+\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (7)
@@
TABLE VI: Fitted parameter values for law 7
@@
$L(N,D,0)=e(1-0)^{\gamma}+\left(a(1-0)^{\alpha}+c\cdot 0\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}.$ (8)
@@ V Conclusions
Current large language models result in higher accuracies and exhibit few-shot learning and even human-like emergent abilities on a wide range of language tasks. However, these models are costly to train. As LLMs get larger and more expensive, it is crucial for researchers and developers to estimate the expected performance beforehand. Scaling laws thus are critical. But while architecture-specific scaling laws give a fine-grained estimation, as more and more new architecture and training methods are being developed, using different scaling laws for each variant becomes impractical. Our generalized scaling law covers densely activated, pruned, and MoE models. Empirical results show that our generalized approach can effectively capture the scaling behavior of the architectures. For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with a given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt.

### v2

@@ I Introduction
In recent years, transformer architectures [1] have revolutionized the deep learning approach. These architectures are now the foundation for the majority of popular large language models (LLMs). As significantly increasing the model size, dataset size and the amount of compute used for training, LLM performance improves dramatically. Because of the large size of LLM models and the large amount of data needed for training, however, the overall training of LLMs becomes time-consuming and expensive. As a result, determining the optimal allocation of resources, model size, and training data and estimating the training performance before conducting the actual training are crucial and have become an important topic in artificial intelligence research. Several approaches have been explored to study the scaling behavior of transformer models to propose scaling laws to determine the optimal allocation of compute, to select optimal model size and data volume for training foundation models efficiently, and to estimate the performance prior to training. Until now, however, nearly all the approaches [2, 3, 4, 5, 6, 7, 8] have been architecture specific (dense or sparse). In this work, we generalize the existing empirical scaling laws and propose a generalized scaling law for both dense and sparse models that enables accurate performance prediction and optimal resource allocation across different architectures.
@@
Scaling laws—empirical relationships often expressed as power laws—describe a model’s performance based on allocated resources. These laws help determine the optimal model size and data volume for a given compute budget to achieve optimal pretraining performance. Several such empirical scaling laws based on different LLM architectures have been proposed in recent years [2, 3, 4, 5, 6, 7, 8, 29]. Others are based on training precision [30], distillation [31], and quantization [32]. Several studies also have been done on vision [4], audio [33, 34], and mixed-modal language models [35]. This interest in scaling laws for different domains makes them pivotal in deep neural network research.
@@ II-A Dense Scaling Laws
Traditional transformer models are dense in nature. Kaplan et al. [2] studied the behavior of LLMs with a series of experiments and different hyperparameters: model size, model shape, total compute, data volume, and batch size. The authors observed that performance or model loss during pretraining follows a power law relationship with the total number of parameters, the total number of training tokens, and the total compute budget. They established that model loss depends strongly on scale, which consists of the number of model parameters $N$, the dataset size (number of tokens) $D$, and the amount of compute $C$ used for training. It also depends mildly on other architectural hyperparameters such as model shape: depth and width. For dense models, the compute $C$ (training FLOPs) required to train a transformer model with $N$ parameters and $D$ tokens is $C=6ND$ [2] because approximately 6 floating point operations are needed for every parameter in the model and every training sample. They suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times to avoid overfitting. They represented training loss as a function of number of parameters and data size and proposed the following empirical scaling law:
$L(N,D)=\left[\left(\frac{N_{C}}{N}\right)^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{C}}{D}\right]^{\alpha_{D}}$ (1)
@@
Using three different approaches—fixed model size with varying numbers of training tokens, IsoFLOP profiling, and parametric function fitting—Hoffmann et al.  [3] extended the work in [2] and concluded that model size and total number of tokens should be increased in proportion to avoid overfitting. In other words, if the total number of parameters or model size is doubled, data volume or total number of tokens should be doubled as well. They represented training loss as a combination of irreducible loss or entropy $e$, a parameter-dependent term or the loss incurred by a model of fixed parameter size, and a data size term or the loss induced by a fixed number of training steps. Based on their observations, the authors proposed the following scaling law:
@@ II-B1 Pruned Transformer Models
$L(N,D,S)=e+\left(a_{S}(1-S)^{b_{S}}+c_{S}\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (4)
@@ II-B2 MoE Models
$L(N,D,S)=e+\frac{a}{N^{\alpha}}+\frac{b}{D^{\beta}}+\frac{c}{(1-S)^{\lambda}}+\frac{d}{(1-S)^{\delta}N^{\gamma}}$ (5)
@@ III Generalizing Scaling Laws
Where $0\leq S<1$ because the active parameters in a model cannot be zero. When $S=0$, it means that the total parameters is the same as the active parameters. Therefore, $\frac{N}{1-S}$ is the total number of parameters. In the parameter-dependent second term in Equation 2, we replace $N$ with $\frac{N}{1-S}$; the parameter-dependent term thus becomes $\frac{a}{\left({\frac{N}{1-S}}\right)^{\alpha}}$, which is simplified as $\frac{a(1-S)^{\alpha}}{N^{\alpha}}$. Similar to the approach in [4] we introduce an upper-bound constant sparsity factor $c$, which is caused by the sparsity; but this factor, which contributes to the loss, depends on the sparsity $S$. So we define the parameter-dependent term as $\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}$. For sparsity $S=0$ or dense models, this term is the same as the original parameter-dependent term in Equation 2.
We have noticed that different approaches adjusted the entropy $e$ in different ways, as shown in Tables II, IV, and V. The value of $e$ is impacted by the sparsity. We adjust the value of $e$ by multiplying it with $(1-S)^{\gamma}$, where $\gamma$ is a constant. Accumulating these changes, we propose the following scaling law:
$L(N,D,S)=e(1-S)^{\gamma}+\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (7)
@@
TABLE VI: Fitted parameter values for Equation 7
@@
$L(N,D,0)=e(1-0)^{\gamma}+\left(a(1-0)^{\alpha}+c\cdot 0\right)\frac{1}{N^{\alpha}}+\frac{b}{D^{\beta}}$ (8)
@@ V Conclusions
Current LLMs result in higher accuracies and exhibit few-shot learning and even human-like emergent abilities on a wide range of language tasks. However, these models are costly to train. As LLMs get larger and more expensive, it is crucial for researchers and developers to estimate the expected performance beforehand. Scaling laws thus are critical. But while architecture-specific scaling laws give a fine-grained estimation, as more and more new architecture and training methods are being developed, using different scaling laws for each variant becomes impractical. Our generalized scaling law covers densely activated, pruned, and MoE models. Empirical results show that our generalized approach can effectively capture the scaling behavior of the architectures. For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with a given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt.